OpenStudy (anonymous):

Add and answer in simplest fractional form. 3/8 + -1/8

OpenStudy (angela210793):

3/8+1/8=4/8=1/2 3/8-1/8=2/8=1/4

OpenStudy (anonymous):

1/2

OpenStudy (stormfire1):

Angela's answer is for \[3/8 \pm 1/8\]If the problem is really \[3/8 + (-1/8)\]then her second answer is the only answer.

OpenStudy (angela210793):

really????????????? tht's not my fault :/

OpenStudy (stormfire1):

No one said there was a problem with the answers you posted. I just wanted to clarify that it really depends on what her original question was...suspecting that it might have been 3/8 + (-1/8) :)